# Unified World Models: Coupling Video and Action Diffusion for Pretraining on Large Robotic Datasets

### v2

@@ III-C Architecture
We model the UWM as a diffusion transformer shown in Fig. 2 and 3. The model predicts actions and observation noises $\epsilon_{a}$ and $\epsilon_{o^{\prime}}$ given current observations $o$, noisy actions $a_{t_{a}}$, noisy observations $o^{\prime}_{t_{o^{\prime}}}$, action timestep $t_{a}$, and observation timestep $t_{o}$. The actions are action chunks of length $h_{a}$. The current observations $o$ and next observations $o^{\prime}$ are frame-stacked observations of length $h_{o}$ from $n_{c}$ camera views.
@@ III-D Training Paradigms
In this work, we evaluate the effectiveness of UWM as a pretraining method for learning the dynamics information from large multitask robotic datasets. To train a UWM, we sample sequences of observations and actions from the dataset, construct $(o,a,o^{\prime})$ tuples, sample random diffusion timesteps $t_{a},t_{o^{\prime}}\sim\mathcal{U}(0,T)$ and optimize the denoising score matching objective in Eq. 1.
Moreover, UWM naturally enables co-training on action-free video data by using diffusion timesteps for masking. Given action-free video samples, instead of sampling the action timestep randomly, we fix the action timestep to $T$ and fill in the missing actions with random noise $\epsilon_{a}\in\mathcal{N}(0,1)$ and optimize the same loss in Eq. 1. We validate the effectiveness of co-training on videos in our experiments in Section. IV-B.
@@ IV Experiments
Figure 4: Visualization of datasets used for pretraining and finetuning. The pretraining/cotraining dataset consists of diverse tasks performed by Franka robots in various environments to ensure broad generalization capabilities. The finetuning datasets include five tasks, each designed to evaluate task-specific performance under controlled conditions.
@@ IV-A Baselines
    Diffusion Policy (DP) [11] is a behavior cloning method that fits a conditional diffusion model to a dataset of expert observation-action data. While the original framework is evaluated in single-task settings, we extend it to the pretraining-finetuning setting by fitting a model to the behavior distribution of a multitask dataset and then finetuning it to the downstream task demonstrations. We compare to DP as a baseline to validate the effectiveness of the additional supervisory signals in UWM. To minimize the discrepancy from UWM, we adopt a diffusion transformer backbone similar to [15] instead of the original UNet architecture [10]. Unlike the encoder-decoder architecture used in [15], however, we only employ the decoder side of the architecture.
@@
    PAD [19] is a joint video-action diffusion model that learns a joint distribution of actions and future observations conditioned on current observations. The key conceptual difference between PAD and UWM is the decoupling of timesteps between actions and next-observations. In addition, PAD conditions the model on the current observations by concatenating the clean latents of the current observations to the noisy latents of the next observations along the channel dimension, similar to Stable Video Diffusion [7]. This is different from the AdaLN condition in UWM. PAD supports co-training on videos by masking the action tokens with a learned mask token.
@@
    GR1 [45] is a video-action transformer model that predicts actions and future image observations conditioned on current image observations. Contrary to other baselines, GR1 does not model a distribution over data using a diffusion generation process. Instead, it directly regresses the actions and images by minimizing a least squares loss. We compare to GR1 to validate the effectiveness of diffusion as a pretraining objective relative to supervised regression. GR1 supports co-training on videos by masking the action tokens with a learned mask token.
IV-B Real Robot Experiments
IV-B1 Setup
To evaluate UWM and baselines as pretraining methods, we leverage the DROID dataset [25] as a source of pretraining data. The DROID dataset is a diverse dataset consisting of robot trajectories collected across various institutions and operators, covering a large variety of tasks, camera positions and backgrounds in natural settings. We curate a pretraining dataset by sampling a subset of 2000 trajectories from the DROID dataset based on location (Fig 4, top row). For methods that support co-training on videos (e.g. GR-1, PAD, and UWM), we want to additionally evaluate their capability of learning from action-free videos. To this end, we curate another 2000 trajectories from the rest of the dataset to use as videos (Fig 4, middle row).
Figure 5: Real-world setup for real robot tasks: Stack-Bowls, Block-Cabinet, Paper-Towel, Hang-Towel, and Rice-Cooker. The first row illustrates the initial configurations for each task, while the second row demonstrates successful task completions. The third row highlights the out-of-distribution (OOD) configurations designed to evaluate the robustness of each method.
  • Stack-Bowls aims to train robotic controllers to place the pink bowl into the blue bowl across various positions.
  • Block-Cabinet aims to open the cabinet and grasp a small red block from the table to place it in the cabinet.
  • Paper-Towel involves precisely grasping a paper towel from a cabinet and placing it upright on a wooden stand on the table.
  • Rice-Cooker (long horizon) involves pouring a cup of rice into the inner bin of a rice cooker, and placing the inner bin on the rice cooker.
We pretrain all methods on the pretraining / co-training datasets for 100K steps and then finetune to the above-mentioned evaluation tasks (task-specific parameters shown in Table. VII.) Specifically, for cotraining experiments, we mix up the robot and video datasets and sample batches uniformly from the mixture dataset, where each batch may contain action-labeled and action-free data. We then apply the method-specific masking techniques and optimize the cotraining loss. For each task, we evaluate in scenarios approximately similar to those encountered during data collection (referred to as in-distribution), and we also construct an out-of-distribution evaluation setting by introducing distractions that are unseen in the finetuning dataset, as shown in Fig 5. To ensure statistically significant evaluation, we test each task on a fixed set of randomly chosen initialization positions. We provide details for the task-specific setups in Appendix -C1.
IV-B2 Discussion
@@
We report the results on the real robot experiments across three tasks in Table I and the average performance in Figure 6. For each method and each task, we provide results in the in-distribution (ID) and out-of-distribution (OOD) scenarios. Furthermore, for methods that support co-training on videos, we additionally report the finetuning results of co-trained model (separated by ”/”).
We first examine the pretraining results on the in-distribution setting. This set of experiments reflect the models’ ability to accurately capture the expert policy’s distribution. We find that UWM achieves the highest success rates across all five tasks among the methods, surpassing the best baseline by as much as 20%. This demonstrates the strength of coupled action-video diffusion in absorbing rich dynamic information from multitask datasets. In particular, since the model is trained to capture all possible conditional and marginal distributions, it is instilled with an understanding of the causal relationship between actions and image observations, explaining its superior performance compared to joint prediction models such as GR1 and PAD. We observe that GR1 consistently outputs the second best results, establishing a strong baseline performance for deterministic regressive models. On the other hand, diffusion policies fail to leverage the rich and dynamic pixel information in the pretraining datasets, being inefficient at learning from diverse multitask trajectories. Finally, due to a combination of conditioning methods and joint image-action diffusion, PAD achieves the lowest success across the board. We attribute its low performance largely to the conditioning via concatenation. Compared to AdaLN in UWM and DP which takes in image features preprocessed by an encoder, PAD takes in raw pixels, thus needing to incorporate the feature extraction in the same transformer model. This limits its performance at accurately capturing the conditional action distribution without expanding model capacity.
IV-C Simulated Experiments
@@
Although our method showed an improvement over baselines, we note that the improvement in OOD scenarios is much less than their real world counterparts. We hypothesize this to be an artifact of current simulations having simpler dynamics than what we see in the real world.
@@ IV-D1 Forward Dynamics
Figure 8: Visualization of the forward dynamics model predictions. The model accurately predicts the robot and object poses conditioned on the initial observation and actions.
@@ IV-D2 Inverse Dynamics
We evaluate the inverse dynamics mode of UWM on trajectory tracking, where we provide a reference expert trajectory and query the inverse dynamics model to track it. Specifically, for each reference trajectory, we reset the simulation environment to match the exact initial state of this trajectory. At each step, we take the ground truth future observations from the trajectory and use the inverse dynamics mode of a finetuned UWM to generate corresponding actions. Table III shows the results of tracking 50 trajectories from the LIBERO training datasets. We find that given the same time limit as the trajectory length, the inverse dynamics model achieves a higher success rate than the policy. This result indicates that actions generated by the inverse dynamics adhere more closely to the reference trajectory. We note that while the policies deviate from the reference trajectories, they eventually recover and solve the tasks given enough time.
@@ IV-D4 Real-World Learning from Scratch
Figure 10: Training models from scratch vs finetuning pretrained models. UWM scales more effectively from pretraining than DP.
IV-D5 Ablation of Learning Objectives
To evaluate whether the performance gain of UWM is a result of dynamics prediction or pure reconstruction, we pretrain a UWM to reconstruct the current observations instead of the future observations. This incentivizes the model to learn about image features, but not about temporal dynamics. Table. V shows that while reconstructing the current observations improves upon the base DP architecture with no image reconstruction, we find it advantageous to reconstruct future observations. This indicates that our model benefits from predicting dynamics rather than purely just image features.
TABLE V: Ablation of Learning Objectives
Stack-Bowls Block-Cabinet
UWM Reconstruct Future Obs 0.86 0.76
UWM Reconstruct Current Obs 0.70 0.66
DP (No Reconstruction) 0.48 0.60
@@ VI Discussion
In this work, we present Unified World Models, a diffusion based framework that naturally unifies policy learning and world modeling into a single flexible framework. We instantiate UWM with a coupled conditional diffusion process using separate timesteps for actions and future observations. During training, the model is exposed to all combinations of timesteps covering various conditional and marginal distributions, instilling the model with a understanding of the causal relationship between actions and future observations. This distinguishes UWM from traditional imitation learning approaches, which often lack a nuanced understanding of causal dependencies. Moreover, the independent diffusion timesteps allow for a natural connection between noising and partial masking, enabling the use of action-free videos for co-training, as well as for marginalization and conditioning of the variables by appropriately setting timesteps. The resulting model is then able to flexibly perform inference as a policy, a video prediction model, a forward dynamics model, and an inverse dynamics model. We show through a thorough experimental evaluation that UWM provide significant gains over imitation learning across the board by enhancing large scale pretraining from robotic datasets.
@@ A-A1 Ablations of Design Choices
TABLE VIII: Ablation of design choices
@@
Results in Table VIII show that adding registers to the transformer help improve the model performance. We hypothesize that adding registers facilitate the exchange of information between actions and latent image patches, which are distinct modalities. We also found that replacing AdaLN conditioning with cross attention results in worse performance. One possible explanation is that action prediction tasks benefit more from AdaLN’s global modulation than from the per-token local modulation provided by cross-attention. We note that this finding may not apply to other modalities such as language.
A-A2 Learning from Internet videos

### v3

@@ III-C Architecture
We model the UWM as a diffusion transformer as shown in Fig. 2 and 3. The model predicts actions and observation noises $\epsilon_{a}$ and $\epsilon_{o^{\prime}}$ given current observations $o$, noisy actions $a_{t_{a}}$, noisy observations $o^{\prime}_{t_{o^{\prime}}}$, action timestep $t_{a}$, and observation timestep $t_{o}$. The actions are action chunks of length $h_{a}$. The current observations $o$ and next observations $o^{\prime}$ are frame-stacked observations of length $h_{o}$ from $n_{c}$ camera views.
@@ III-D Training Paradigms
In this work, we evaluate the effectiveness of UWM as a pretraining method for learning the dynamics information from large multitask robotic datasets. To train a UWM on robot data, we sample sequences of observations and actions from the dataset, construct $(o,a,o^{\prime})$ tuples, sample random diffusion timesteps $t_{a},t_{o^{\prime}}\sim\mathcal{U}(0,T)$, and optimize the denoising score matching objective in Eq. 1.
Moreover, UWM naturally enables co-training on action-free video data by using diffusion timesteps for masking. Given action-free video samples, instead of sampling the action timestep randomly, we fix the action timestep to $T$, impute the missing actions with random noise $\epsilon_{a}\in\mathcal{N}(0,1)$, optimize the same loss in Eq. 1. We validate the effectiveness of co-training on videos in our experiments in Section. IV-B.
@@ IV Experiments
Figure 4: Visualization of datasets used for pretraining and finetuning. The pretraining and cotraining dataset consists of diverse tasks performed by Franka robots in various environments to ensure broad generalization capabilities. The finetuning datasets include five tasks, each designed to evaluate task-specific performance under controlled conditions.
Figure 5: Setup for real robot tasks: Stack-Bowls, Block-Cabinet, Paper-Towel, Hang-Towel, and Rice-Cooker. The first row illustrates the initial configurations for each task. The second row demonstrates successful task completions. The third row highlights the out-of-distribution (OOD) configurations designed to evaluate the robustness of each method.
@@ IV-A Baselines
    Diffusion Policy (DP) [11] is a behavior cloning method that fits a conditional diffusion model to a dataset of expert observation-action data. We extend the framework to the pretraining-finetuning setting by fitting a model to the behavior distribution of a multitask dataset and then finetuning it to the task-specific demonstrations. We compare to DP as a baseline to validate the effectiveness of the additional supervisory signals in UWM. To minimize the discrepancy from UWM, we adopt a diffusion transformer backbone similar to [15] instead of the original UNet architecture [10].
@@
    PAD [19] is a video-action diffusion model that learns a joint distribution of actions and future observations conditioned on current observations. The key conceptual difference between PAD and UWM is the decoupling of timesteps between actions and next-observations. In addition, PAD conditions the model on the current observations by concatenating the clean latents of the current observations to the noisy latents of the next observations along the channel dimension, similar to [7]. PAD supports co-training on videos by masking the action tokens with a learned mask token.
@@
    GR1 [45] is a video-action transformer model that predicts actions and future image observations conditioned on current image observations. Unlike other baselines, GR1 does not model a distribution over data using a diffusion process. Instead, it directly regresses the actions and images by minimizing a least squares loss. We compare to GR1 to validate the effectiveness of diffusion as a pretraining objective relative to regression. GR1 supports co-training on videos by masking the action tokens with a learned mask token.
@@
IV-B Real Robot Experiments
IV-B1 Setup
To evaluate UWM and baselines as pretraining methods, we leverage the DROID dataset [25] as a source of pretraining data. The DROID dataset is a diverse dataset consisting of robot trajectories collected across various institutions and operators, covering a large variety of tasks, camera positions and backgrounds in natural settings. We curate a pretraining dataset by sampling a subset of 2000 trajectories from the DROID dataset based on location (Fig 4, top row). For methods that support co-training on video data (e.g. GR-1, PAD, and UWM), we additionally evaluate their capability of learning from action-free videos. To this end, we sample another 2000 trajectories from the rest of the DROID dataset and remove their action annotations to use as videos (Fig 4, middle row).
  • Stack-Bowls involves picking up the pink bowl and stacking it on top of the blue bowl.
  • Block-Cabinet involves opening the cabinet, grasping a small red block from the table, and placing it in the cabinet.
  • Paper-Towel involves precisely grasping a paper towel roll from the cabinet and placing it upright on a wooden stand on the table.
  • Rice-Cooker (long horizon) involves pouring a cup of rice into the inner pot of a rice cooker, and putting the inner pot in the rice cooker.
We train all methods on the pretraining / co-training datasets for 100K steps and then finetune to the evaluation tasks (task-specific parameters shown in Table. VI.) For cotraining experiments, we mix up the robot and video datasets and sample batches uniformly from the mixture dataset, where each batch may contain action-labeled and action-free data. We then apply the method-specific masking techniques and optimize the cotraining loss. For each task, we evaluate in scenarios approximately similar to those encountered during data collection (referred to as in-distribution), and we also construct an out-of-distribution evaluation setting by introducing distractions that are unseen in the finetuning dataset, as shown in Fig 5. To ensure statistically significant evaluation, we test each task on a fixed set of randomly chosen initialization positions. We provide details for the task-specific setups in Appendix -C1.
IV-B2 Discussion
@@
We report the results on the real robot experiments in Table I and the average performance in Figure 6. For each method an task, we provide results in the in-distribution (ID) and out-of-distribution (OOD) scenarios. Furthermore, for methods that support co-training on videos, we additionally report the results of co-trained models (separated by ”/”).
We first examine the pretraining results in the in-distribution setting. This set of experiments reflect the models’ ability to accurately capture the expert policy’s distribution. We find that UWM achieves the highest success rates across all five tasks among the methods, surpassing the best baseline by as much as 20%. This demonstrates the strength of coupled action-video diffusion in absorbing rich dynamic information from multitask datasets. In particular, since the model is trained to capture all possible conditional and marginal distributions, it is instilled with an understanding of the causal relationship between actions and image observations, explaining its superior performance compared to joint prediction models such as GR1 and PAD. GR1 consistently outputs the second best results, establishing a strong baseline performance for deterministic regressive models. Diffusion Policies fail to leverage the rich and dynamic pixel information in the pretraining datasets, being inefficient at learning from diverse multitask trajectories. PAD achieves the lowest success across the board. We attribute its low performance largely to the conditioning via concatenation. Compared to UWM which takes in image features preprocessed by an encoder, PAD takes in raw pixels, thus needing to incorporate the feature extraction in the same transformer model. This limits its performance at accurately capturing the conditional action distribution without expanding model capacity.
IV-C Simulated Experiments
@@
Although our method showed an improvement over baselines, we note that the improvement in OOD scenarios is less than the real world experiments. We hypothesize this to be an artifact of current simulations having simpler dynamics than what we see in the real world.
@@ IV-D1 Forward Dynamics
Figure 8: Visualization of the forward dynamics predictions. The model accurately predicts the robot and object poses conditioned on the initial observation and actions.
@@ IV-D2 Inverse Dynamics
We evaluate the inverse dynamics mode of UWM on trajectory tracking, where we provide a reference expert trajectory and query the inverse dynamics model to track it. Specifically, for each reference trajectory, we reset the simulation environment to match the exact initial state of this trajectory. At each step, we take the ground truth future observations from the trajectory and use the inverse dynamics mode of a finetuned UWM to generate corresponding actions. Table III shows the results of tracking 50 trajectories from the LIBERO training datasets. We find that given the same time limit as the trajectory length, the inverse dynamics model achieves a higher success rate than the policy. This implies that actions generated by the inverse dynamics adhere more closely to the reference trajectory. We note that while the policies deviate from the reference trajectories, they eventually recover and solve the tasks given enough time.
@@ IV-D4 Real-World Learning from Scratch
Figure 10: Training models from scratch vs finetuning pretrained models. UWM scales more effectively with pretraining than DP.
@@ VI Discussion
In this work, we present Unified World Models, a diffusion based framework that unifies policy learning and world modeling into a single flexible framework. We instantiate UWM with a coupled conditional diffusion process using separate timesteps for actions and future observations. During training, the model is exposed to all combinations of timesteps covering various conditional and marginal distributions, instilling the model with an understanding of the causal relationship between actions and future observations. This distinguishes UWM from traditional imitation learning approaches, which often lack a nuanced understanding of causal dependencies. Moreover, the independent diffusion timesteps allow for a natural connection between noising and partial masking, enabling the use of action-free videos for co-training, as well as for marginalization and conditioning of the variables by appropriately setting timesteps. The resulting model is then able to flexibly perform inference as a policy, a video prediction model, a forward dynamics model, and an inverse dynamics model. We show through a thorough experimental evaluation that UWM provides significant gains over imitation learning across the board by enhancing large scale pretraining from robotic datasets.
@@ A-A1 Ablations of Design Choices
TABLE VII: Ablation of design choices
@@
Results in Table VII show that adding registers to the transformer help improve the model performance. We hypothesize that adding registers facilitate the exchange of information between actions and latent image patches, which are distinct modalities. We also found that replacing AdaLN conditioning with cross attention results in worse performance. One possible explanation is that action prediction tasks benefit more from AdaLN’s global modulation than from the per-token local modulation provided by cross-attention. We note that this finding may not apply to other modalities such as language.
A-A2 Ablation of Learning Objectives
To evaluate whether the performance gain of UWM is a result of dynamics prediction or pure reconstruction, we pretrain a UWM to reconstruct the current observations instead of the future observations. This incentivizes the model to learn about image features, but not about temporal dynamics. Table. VIII shows that while reconstructing the current observations improves upon the base DP architecture with no image reconstruction, we find it advantageous to reconstruct future observations. This indicates that our model benefits from predicting dynamics rather than purely just image features.
TABLE VIII: Ablation of Learning Objectives
Stack-Bowls Block-Cabinet
UWM Reconstruct Future Obs 0.86 0.76
UWM Reconstruct Current Obs 0.70 0.66
DP (No Reconstruction) 0.48 0.60
A-A3 Learning from Internet videos